5. A graph of the function fx= x3-3x2 shown below. Each of the following graphs (a-d) show a function transformed from this one. Find a formula for each function. Explain how the function was changed. (1.25 pts)

- N -2-) E- . آشناير 12 (4.23
Answer #1

For 1st graph :

When compared to the graph of f(x) = x​​​​​​3​​​​​ - 3x2, it is shifted 1 unit towards left i.e.towards negative side.

So, k(x) = (x+1)3- 3(x+1)2 = x​​​​​​3 - 3x - 2

For 2nd graph :

When compared to f(x), graph has shifted towards positive y axis by 1 unit and positive x axis by 2 units.

So, y - 1 = (x-2)3 - 3(x-2)2 = x​​​​​​3 - 9x2 + 24x -20

t(x) = y = x​​​​​​3 - 9x2 + 24x - 19